Find the orthogonal trajectories of the given family of curves y= Ce^-mx; with m held fixed
1
Expert's answer
2020-12-21T18:08:00-0500

"y = Ce^{-mx}"


First, this family of curves satisfies the differential equation


"\begin{aligned}y = Ce^{-mx} \quad \implies \quad y' = -mCe^{-mx} \end{aligned}"


Since "C = ye^{mx}" we then have


"y' = -my"


Letting "f(x,y) = -my", we know the orthogonal trajectories are the curves which satisfy a differential equation


"y' = \dfrac{-1}{f(x,y)} \implies y' = \dfrac{1}{my}"


Hence, the orthogonal trajectories that satisfy the differential equation are the curves


"\begin{aligned}
y' = \frac{1}{my} \qquad \implies \quad & myy' = 1\\
\implies \quad & m\int{ydy} = dx\\
\implies \quad & \dfrac{my^2}{2} = x + C\\
\implies \quad & my^2 - 2x = 2C
\end{aligned}"

where C is an arbitrary constant.


The orthogonal trajectories for the family of straight lines are parabolas (y² = ax + c)